# It’s All About Your Sketch: Democratising Sketch Control in Diffusion Models

### 12 Mar 2024 v1

@@ 2 Related Works
Diffusion Model for Vision Tasks.   Diffusion models [24, 23, 73] have now become the gold-standard for different controllable image generation frameworks like DALL-E [56], Imagen [63], T2I-Adapter [54], ControlNet [90], etc. Besides image generation, several methods like Dreambooth [62], Imagic [31], Prompt-to-Prompt [21], SDEdit [51], SKED [53] extend it for realistic image editing. Beyond image generation and editing, diffusion model is also used in several downstream vision tasks like recognition [42], semantic [2] and panoptic [84] segmentation, image-to-image translation [78], medical imaging [14], image correspondence [77], image retrieval [38], etc.
Sketch for Visual Content Creation.  Following its success in sketch-based image retrieval (SBIR) [65, 10, 3], sketches are now being used in other downstream tasks like saliency detection [6], augmented reality [49, 50], medical image analysis [34], object detection [13], class-incremental learning [4], etc. Apart from the plethora of sketch-based $2D$ and $3D$ image generation and editing frameworks [35, 46, 59, 20, 90, 54, 80, 53, 81], sketches are also getting significant traction in other visual content creation tasks like animation generation [72] and inbetweening [71], creative garment design [45, 11], manga restoration [82], CAD modelling [43, 88], anime editing [27], etc.
@@ 3 Revisiting Diffusion Model (DM)
$\mathcal{L}_{\mathrm{SD}}=\mathbb{E}_{\mathbf{z}_{t},t,{d},\epsilon}({||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{T}(d))||}_{2}^{2})$ (2)
During inference, SD discards $\mathcal{E}(\cdot)$, directly sampling a noisy latent $\mathbf{z}_{T}$ from a Gaussian distribution [60]. It then estimates noise from $\mathbf{z}_{T}$ iteratively for $T$ iterations via $\epsilon_{\theta}$ (conditioned on $d$) to obtain a clean latent $\hat{\mathbf{z}}_{0}$. The frozen decoder generates the final image as $\hat{\mathbf{x}}_{0}=\mathcal{D}(\hat{\mathbf{z}}_{0})$ [60].
@@ 5.1 Sketch Adapter
$\mathcal{L}_{\mathrm{SD}}=\mathbb{E}_{\mathbf{z}_{t},t,{s},\epsilon}({||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))||}_{2}^{2})\vspace{-0.1cm}$ (3)
@@ 5.3 Super-concept Preservation Loss
$\mathcal{L}_{\mathrm{reg}}=||\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{T}(d))-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))||_{2}^{2}\vspace{-0.1cm}$ (6)
@@ 5.4 Abstraction-aware Importance Sampling
where, $\mathcal{S}_{\omega}(\cdot)$ is our abstraction-aware $t$-sampling function, where increasing or decreasing $\omega\in(0,1]$, controls the skewness of this sampling probability density function. Pushing $\omega$ towards $1$ increases the probability of sampling a larger $t$ value. We aim to make this skewness-controlling $\omega$ value sketch-abstraction specific (Fig. 4).
@@
$\footnotesize\mathcal{L}_{\mathrm{abs}}=-\mathrm{log}\frac{e^{s~{}\mathrm{cos}(\theta_{y_{i}}+m(\mathbf{s}_{i}))}}{e^{s~{}\mathrm{cos}(\theta_{y_{i}}+m(\mathbf{s}_{i}))}+\sum_{j\neq y_{i}}e^{s~{}\mathrm{cos}~{}\theta_{j}}}+\lambda_{g}g(\mathbf{s}_{i})$ (8)
@@
In summary, we train the sketch adapter $\mathcal{A}(\cdot)$ using sketch-abstraction-aware $t$-sampling with a total loss of $\mathcal{L}_{\mathrm{total}}$=$\lambda_{1}\mathcal{L}_{\mathrm{SD}}$+$\lambda_{2}\mathcal{L}_{\mathrm{SBIR}}$+$\lambda_{3}\mathcal{L}_{\mathrm{reg}}$. During inference, we compute the abstraction score of the input sketch, taking $l_{2}$-norm of classifier feature. Based on the abstraction level, we perform $t$-sampling. The input sketch passing through $\mathcal{A}$ controls the diffusion procedure and generates the output.
@@ 6.2 Ablation on Design
[ii] Why Discriminative Learning? Fine-grained discriminative loss (Eq. 5) helps the conditioning process by distilling knowledge learned inside a pre-trained FG-SBIR model. As seen in Tab. 2, a noticeable FGM drop ($44.44\%$) for w/o Discriminative learning indicates that fine-grained sketch-conditioning is incomplete without explicit discriminative learning via $\mathcal{L}_{\mathrm{SBIR}}$.
@@ 7 Conclusion
Subhadeep Koley1,2    Ayan Kumar Bhunia1    Deeptanshu Sekhri1    Aneeshan Sain1,2
Pinaki Nath Chowdhury1,2    Tao Xiang1,2    Yi-Zhe Song1,2

### 20 Mar 2024 v2

@@ 2 Related Works
Diffusion Models for Vision Tasks.   Diffusion models [25, 24, 74] have now become the gold-standard for different controllable image generation frameworks like DALL-E [57], Imagen [64], T2I-Adapter [55], ControlNet [90], etc. Besides image generation, several methods like Dreambooth [63], Imagic [32], Prompt-to-Prompt [22], SDEdit [52], SKED [54] extend it for realistic image editing. Beyond image generation and editing, diffusion model is also used in several downstream vision tasks like recognition [43], semantic [2] and panoptic [84] segmentation, image-to-image translation [79], medical imaging [15], image correspondence [78], image retrieval [39], etc.
Sketch for Visual Content Creation.  Following its success in sketch-based image retrieval (SBIR) [66, 11, 3], sketches are now being used in other downstream tasks like saliency detection [6], augmented reality [50, 51], medical image analysis [35], object detection [14], class-incremental learning [4], etc. Apart from the plethora of sketch-based $2D$ and $3D$ image generation and editing frameworks [36, 47, 60, 21, 90, 55, 81, 54, 82], sketches are also getting significant traction in other visual content creation tasks like animation generation [73] and inbetweening [72], garment design [46, 12], caricature generation [10], CAD modelling [44, 88], anime editing [28], etc.
@@ 3 Revisiting Diffusion Model (DM)
$\mathcal{L}_{\text{SD}}=\mathbb{E}_{\mathbf{z}_{t},t,{d},\epsilon}({||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{T}(d))||}_{2}^{2})$ (2)
During inference, SD discards $\mathcal{E}(\cdot)$, directly sampling a noisy latent $\mathbf{z}_{T}$ from a Gaussian distribution [61]. It then estimates noise from $\mathbf{z}_{T}$ iteratively for $T$ iterations via $\epsilon_{\theta}$ (conditioned on $d$) to obtain a clean latent $\hat{\mathbf{z}}_{0}$. The frozen decoder generates the final image as: $\hat{\mathbf{x}}_{0}=\mathcal{D}(\hat{\mathbf{z}}_{0})$ [61].
@@ 5.1 Sketch Adapter
$\mathcal{L}_{\text{SD}}=\mathbb{E}_{\mathbf{z}_{t},t,{s},\epsilon}({||\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))||}_{2}^{2})\vspace{-0.1cm}$ (3)
@@ 5.3 Super-concept Preservation Loss
$\mathcal{L}_{\text{reg}}=||\epsilon_{\theta}(\mathbf{z}_{t},t,\mathbf{T}(d))-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{A}(\mathbf{V}(s)))||_{2}^{2}\vspace{-0.1cm}$ (6)
@@ 5.4 Abstraction-aware Importance Sampling
where, $\mathcal{S}_{\omega}(\cdot)$ is our abstraction-aware $t$-sampling function, where increasing or decreasing $\omega\in(0,1]$, controls the skewness of this sampling probability density function. Pushing $\omega$ towards $1$ increases the probability of sampling a larger $t$ value (Fig. 4). We aim to make this skewness-controlling $\omega$ value sketch-abstraction specific.
@@
$\footnotesize\mathcal{L}_{\text{abs}}=-\mathrm{log}\frac{e^{s~{}\mathrm{cos}(\theta_{y_{i}}+m(\mathbf{s}_{i}))}}{e^{s~{}\mathrm{cos}(\theta_{y_{i}}+m(\mathbf{s}_{i}))}+\sum_{j\neq y_{i}}e^{s~{}\mathrm{cos}~{}\theta_{j}}}+\lambda_{g}g(\mathbf{s}_{i})$ (8)
@@
In summary, we train the sketch adapter $\mathcal{A}(\cdot)$ using sketch-abstraction-aware $t$-sampling with a total loss of $\mathcal{L}_{\text{total}}$=$\lambda_{1}\mathcal{L}_{\text{SD}}$+$\lambda_{2}\mathcal{L}_{\text{SBIR}}$+$\lambda_{3}\mathcal{L}_{\text{reg}}$. During inference, we compute the abstraction score of the input sketch, taking $l_{2}$-norm of classifier feature. Based on the abstraction level, we perform $t$-sampling. The input sketch passing through $\mathcal{A}$ controls the diffusion procedure and generates the output.
@@ 6.2 Ablation on Design
[ii] Why Discriminative Learning? Fine-grained discriminative loss (Eq. 5) helps the conditioning process by distilling knowledge learned inside a pre-trained FG-SBIR model. As seen in Tab. 2, a noticeable FGM drop ($44.44\%$) for w/o Discriminative learning indicates that fine-grained sketch-conditioning is incomplete without explicit discriminative learning via $\mathcal{L}_{\text{SBIR}}$.
@@ 7 Conclusion
Subhadeep Koley1,2    Ayan Kumar Bhunia1    Deeptanshu Sekhri1    Aneeshan Sain1
Pinaki Nath Chowdhury1    Tao Xiang1,2    Yi-Zhe Song1,2